# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

### v1

@@ II.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs can be measured. This is done by assuming that only the nearest neighbors can affect the accumulation under a gate via crosstalk when pinching off the channels. Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend, eventually cutting off the current in the respective channels. An idealized representation of such a curve is visualized in the left panel in the second row in Fig. 2(e).
@@ III Experimental Demonstration
We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig. 1(a). The device’s schematic consisting of 24 gates is shown in Fig. 3(a). The first layer of gates is made up of four screening gates, Si for , shown in Fig. 3(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in Fig. 3(a) in blue, while the plungers are shown in red. The final layer consists of nine barrier gates, Bi for , shown in Fig. 3(a) in gray.
The goal of BATIS is to form three current channels in the quad-QD device, as indicated with white arrows in Fig. 3(a). BATIS brings a grounded but not characterized device from Fig. 1(a) to a fully operational device in Fig. 1(d). In what follows, we explain how BATIS performs the quad-QD device bootstrapping, testing, and characterizations.
@@
The leakage test needs to be executed only once per cool-down if under fully automatic control. This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
Global Accumulation. Once no leakage is confirmed, BATIS performs the global turn-on stage. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and . During the first two illuminations, all gates were grounded to reset the device after cooldown and verify that no instruments were erratic. After that, for the third illumination, BATIS chose to apply voltages to all screening and reservoir gates. The voltages required to achieve turn-ons at around , calculated following the protocol in Ref. [26], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3. The central screening gate, S2, was left grounded during the illumination.
| Channel | Run 1 | Run 2 | Run 3 | Run 4 | |
| characteristics | |||||
@@
| Vmax () | |||||
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is determined, following the same strategy as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since the ohmic connections are excluded, the resulting matrix has five fewer indexes. As none of the connections leaked, only measurements following the matrix diagonal were taken.
@@
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [26]. We note that for gate S1, BATIS acquired signal over , i.e., twice as long as for all remaining gates. This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
BATIS sets the reservoir gates two standard deviations above the mean of the phenomenological fit it performed on the signal’s gradient. This is to ensure that reservoirs will consistently provide electron flow through the channel as we do not want the cross-talk effects to cause drastic changes in the electron accumulation. At the same time, if the reservoirs are too accumulated, they can form spurious channels behind the regions of interest. To resolve this, the screening gates are set below their pinch-offs in Fig. 3(b) to prevent current flow through spurious channels 111The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative..
The summary of the pinch-off voltages and the operating point selected by BATIS is presented in Table 2. The resulting configuration, where the device is in the correct operating point to perform channel accumulation, is visualized in the central plot in Fig. 1(c).
@@
Finger Gate Characterization. The last stage of BATIS involves finger gate characterization. This characterization is done individually for all of the finger gates on the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others. This can be seen in the raw data pictured in Fig. 3(d).
To ensure that the right point is consistently found, BATIS combines thresholding and slope extraction for each plot to ensure that the pinch-off it finds is robust to noise and spurious fluctuations that do not exceed the threshold. The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 1(d).
@@ IV Summary and outlook
In this paper, we proposed and demonstrated BATIS: a fully automatic procedure for bootstrapping, automated testing, and characterization of accumulation mode, undoped QD devices. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, and the device limits. BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch off and that the current channels can be formed. Testing and characterizing an unknown QD device is essential for identifying suitable global device states that enable the observation of Coulomb blockade. Importantly, the single button press testing and initialization is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional. This result is particularly important as we envision a hot large-scale spin-based quantum processor. These systems will require qubit initialization, relying on a complete and accurate device characterization. One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a partition of a large qubit processor without going to temperatures.
The flexible structure allows BATIS to be scalable to characterize and initialize any accumulation mode linear array of QD and easily adaptable to other QD platforms. In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessarily supported for these devices. However, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
Previous studies have demonstrated a device-agnostic algorithm for certain aspects of QD device tuning, such as setting topology [32, 17] and charge state classification [33, 22]. However, all these algorithms presuppose that device characterization and foundational tuning steps—determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse tuning regime—have already been performed. BATIS fills this gap, making device characterization easily within reach for linear QD array systems and paving the way toward fully automatic control. When linked together, all proposed subroutines can form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, which will accommodate the growth of QD qubit systems on the path to a scalable quantum computer on this platform.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work supported in part by ARO grant nos. W911NF-23-1-0110 and W911NF-17-1-0274. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
@@
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, E. Acuna, R. S. Noah, J. W. Matten, M. P. Jura, J. A. Wright, M. T. Rakher, and M. G. Borselli, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022), pMID: 34806894.
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, L. E. A. Stehouwer, S. D. Oosterhout, A. Sammak, M. Friesen, G. Scappucci, S. L. de Snoo, M. Rimbach-Russ, F. Borsoi, and M. Veldhorst, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, D. Keith, L. Lampert, F. Luthi, F. A. Mohiyaddin, S. Murcia, R. Nair, R. Nahm, A. Nethwewala, S. Neyens, R. D. Raharjo, C. Rogan, R. Savytskyy, T. F. Watson, J. Ziegler, O. K. Zietz, R. Pillarisetty, N. C. Bishop, S. A. Bojarski, J. Roberts, and J. S. Clarke, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
@@
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, E. J. Connors, J. Corrigan, M. J. Curry, D. Keith, R. Kotlyar, L. F. Lampert, M. T. Mądzik, K. Millard, F. A. Mohiyaddin, S. Pellerano, R. Pillarisetty, M. Ramsey, R. Savytskyy, S. Schaal, G. Zheng, J. Ziegler, N. C. Bishop, S. Bojarski, J. Roberts, and J. S. Clarke, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
@@
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, M. J. Carballido, S. Svab, A. V. Kuhlmann, S. Geyer, F. N. M. Froning, H. Moon, M. A. Osborne, D. Sejdinovic, G. Katsaros, D. M. Zumbühl, G. A. D. Briggs, and N. Ares, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
@@
- Wolfe et al. [2024] M. A. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. G. Lagally, R. McDermott, M. Friesen, S. Kolkowitz, and M. A. Eriksson, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
@@
- Note [2] The derivative of the screening gate pinch-off was fit to a function like . The non-central screening gates were set to be before the mean of the derivative.

### v2

@@ II.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs (VP) can be measured. Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend consistent with Eq. 1, eventually cutting off the current in the respective channels following Eq. (2) with . An idealized representation of such a curve is visualized in the left panel in the second row in Fig. 2(e).
@@ III Experimental Demonstration
We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig. 1(a) with 24 gates. The first layer of gates is made up of four screening gates, Si for , shown in Fig. 1(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in Fig. 1(a) in blue, while the plungers are shown in red. The final layer consists of nine barrier gates, Bi for , shown in Fig. 1(a) in gray.
The goal of BATIS is to form three current channels in the quad-QD device as it learned from the gate layout in the configuration, as indicated with white arrows in Fig. 3(a). BATIS brings a grounded but not characterized device from Fig. 1(a) to a fully operational device in Fig. 1(d). In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterizations.
@@
The leakage test needs to be executed only once per cooldown if under fully automatic control. This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
| Voltages | Run 1 | Run 2 | Run 3 | Run 4 | |
@@
| Vmax () |
Global Accumulation. Next, BATIS performs the global accumulation stage. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and . By design, all gates are grounded during the first two illuminations. The first illumination is intended to completely reset the device while the second is to verify that all instruments work as intended. Because BATIS still did not find the gates in the desired range, it chose to apply voltages to all screening and reservoir gates for the third illumination. The voltages required to achieve turn-ons at around , calculated following the protocol in Ref. [21], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3. The central screening gate, S2, was left grounded during the illumination.
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is measured, following the same procedure as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since the ohmic connections are excluded, the resulting matrix has five fewer indexes. As none of the connections leaked, only measurements following the matrix diagonal were taken.
@@
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [21]. We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long as compared to all other screening gates. This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2. A device in this configuration prepared for the channel accumulation stage is shown in Fig. 1(c).
@@
Finger Gate Characterization. The last stage of BATIS involves finger gate characterization. This characterization is done individually for all of the finger gates on the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others. This can be seen in the raw data pictured in Fig. 3(d). The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 1(d).
@@ IV Summary and outlook
In this paper, we propose and demonstrate BATIS, a fully autonomous procedure for bootstrapping, automated testing, and characterization of accumulation mode undoped QD devices. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, the ohmic biases, and the device limits. BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch-off and that all current channels can be formed. Testing and characterizing an unknown QD device is essential for identifying global device states that are suitable for observing the Coulomb blockade. Importantly, the single-button-press testing and initialization process is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional. This result is particularly important as we envision a hot large-scale spin-based quantum processor. These systems will require qubit initialization, relying on a complete and accurate device characterization. One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a portion of a large qubit processor without going to temperatures.
The flexible structure of BATIS allows for scalable characterization and initialization of any accumulation-mode linear array of QDs. It is also easily adaptable to other QD platforms. In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessary for these devices. On the other hand, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.
Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [33, 17] or charge-state classifications [34, 24]. However, these algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed. BATIS fills this gap, bringing device characterization easily within reach for linear QD arrays and paving the way toward fully automatic control. When linked together, the proposed subroutines form an autonomous system for controlling arbitrary QD devices up to any desired level of characterization, putting QD qubit systems on the path to scalable quantum computing.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work was supported in part by ARO grant nos. W911NF-17-1-0274, W911NF-23-1-0110, and W911NF-24-2-0043. The authors gratefully acknowledge the use of facilities and instrumentation in the UW-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF DMR-2309000) and the University of Wisconsin-Madison. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
@@
- Ha et al. [2022] W. Ha, S. D. Ha, M. D. Choi, Y. Tang, A. E. Schmitz, M. P. Levendorf, K. Lee, J. M. Chappell, T. S. Adams, D. R. Hulbert, et. al, A flexible design platform for si/sige exchange-only qubits with low disorder, Nano Lett. 22, 1443 (2022), pMID: 34806894.
- Wang et al. [2024] C.-A. Wang, V. John, H. Tidjani, C. X. Yu, A. S. Ivlev, C. Déprez, F. van Riggelen-Doelman, B. D. Woods, N. W. Hendrickx, W. I. L. Lawrie, et. al, Operating semiconductor quantum processors with hopping spins, Science 385, 447 (2024).
- George et al. [2024] H. C. George, M. T. Mądzik, E. M. Henry, A. J. Wagner, M. M. Islam, F. Borjans, E. J. Connors, J. Corrigan, M. Curry, M. K. Harper, et. al, 12-spin-qubit arrays fabricated on a 300 mm semiconductor manufacturing line (2024), arXiv:2410.16583 [cond-mat.mes-hall] .
- Note [2] For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable. However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.
@@
- Neyens et al. [2024] S. Neyens, O. K. Zietz, T. F. Watson, F. Luthi, A. Nethwewala, H. C. George, E. Henry, M. Islam, A. J. Wagner, F. Borjans, et. al, Probing single electrons across 300-mm spin qubit wafers, Nature 629, 80 (2024).
@@
- Severin et al. [2024] B. Severin, D. T. Lennon, L. C. Camenzind, F. Vigneau, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, M. De Kruijf, et. al, Cross-architecture tuning of silicon and SiGe-based quantum devices using machine learning, Sci. Rep. 14, 17281 (2024).
- Wolfe et al. [2024] M. A. Wolfe, B. X. Coe, J. S. Edwards, T. J. Kovach, T. McJunkin, B. Harpt, D. Savage, M. G. Lagally, R. McDermott, M. Friesen, et. al, Control of threshold voltages in Si/Si0.7 Ge0.3 quantum devices via optical illumination, Physical Review Applied 22, 034044 (2024).
@@
- Note [3] The parameter is defined as part of the configuration file and can be adjusted if needed.